OpenStudy (anonymous):

In Steinhatchee in July, high tide is at noon. The water level is 5 feet at high tide and 1 foot at low tide. Assuming the next high tide is exactly 12 hours later and the height of the water can be modeled by a cosine curve, find an equation for Steinhatchee's water level in July as a function of time (t).

OpenStudy (anonymous):

A. f(t) = 6 cos pi over 2 t + 2 B. f(t) = 2 cos pi over 2 t + 3 C. f(t) = 2 cos pi over 6 t + 3 D. f(t) = 6 cos pi over 6 t + 2